# A Comprehensive Graph Pooling Benchmark: Effectiveness, Robustness and Generalizability

### v3

@@ 4.3 Robustness Analysis
Table LABEL:noise_node presents the results of the robustness analysis for node-level tasks. From Table LABEL:noise_node, we observe the following: Firstly, random attacks on the graph lead to a decrease in performance on node classification tasks, with different types of attacks causing varying degrees of performance degradation. Randomly adding edges has the most negative impact on performance, while randomly deleting edges has the least impact. Secondly, for larger graphs such as Cora, Citeseer, and Pubmed, KMISPool performs the best, whereas for smaller graphs such as Cornell, Texas, and Wisconsin, ASAPool performs better. Appendix F.1 provides results for the robustness analysis of node-level tasks.
As depicted in As shown in Figure 2, the model’s performance generally declines as the noise intensity increases. It is observed that, at the same level of noise, the impact on accuracy is more pronounced on smaller datasets Cora and CiteSeer, while it is relatively minor on larger dataset Pubmed. Among the three types of noise, although the accuracy of nearly all methods decreases amidst fluctuations, KMISPool and PANPool exhibit the strongest robustness, while COPool performs relatively poorly, despite the fact that most pooling methods show very similar performance.
@@ 4.4 Generalizability Analysis
Table 7 and Table 8 presents the performance of different graph pooling methods under out-of-distribution shifts. For the graph-level datasets D&D and NCI1, we implement two types of distribution shifts. The first type is based on the number of nodes, where the smallest 50% of graphs by node count are used as the training set, and the largest 20% as the test set, with the remainder serving as the validation set (Bevilacqua et al., 2021; Chen et al., 2022). Following the same criteria, the second type of out-of-distribution shifts are generated based on graph density (Chen et al., 2022). For the node-level datasets Cora and Citeseer, the first type of out-of-distribution shift is the top 50% of nodes with the highest degrees as the training set, the bottom 25% with the lowest degrees as the test set, and the remaining nodes as the validation set. The second type is based on closeness centrality (the reciprocal of the sum of the shortest path lengths from a node to all other nodes). We use the 50% of nodes with the lowest closeness as the training set, the 25% with the highest closeness as the test set, and the remaining nodes as the validation set. For further details and more experiments for the generalizability analysis, please refer to the Appendix D.4 and Appendix F.2.
@@ 4.5 Further Analysis
Backbone Analysis. Figure 5 presents the performance of four pooling methods based on GCNConv (Kipf & Welling, 2016b), GATConv (Veličković et al., 2017), SAGEConv (Hamilton et al., 2017), and GraphConv (Morris et al., 2019) on four datasets, NCI1, NCI109, PROTEINS, and PROTEINS_FULL. On average, as the backbone models change, most pooling methods exhibit significant performance fluctuations, and no single backbone model consistently maintains a leading position. Except for the PROTEINS_FULL, the performance of GraphConv is relatively better.
@@ 5 Conclusion
Appendix A Related Work
A.1 Graph Classification and Graph Regression
A.2 Node Classification
Node classification aims to classify each node in a graph based on its attributes and relationship with the other nodes (Xiao et al., 2022; Ju et al., 2024). Node classification has various applications in the real world, including social network analysis (Camacho et al., 2020), knowledge graphs (Ye et al., 2022), bioinformatics (Bhagat et al., 2011) and online commerce services (Yu et al., 2023). Graph neural networks have been widely utilized to solve the problem by learning semantics information across nodes by neighborhood propagation. Since graph pooling would reduce the number of nodes, recent works utilize a U-Net architecture (Gao & Ji, 2019), which involves down-sampling and up-sampling with residue connections. In this work, we systematically evaluate the performance of graph pooling on node classification and observe that graph pooling has limited improvement compared with basic graph neural network architectures.
A.3 Previous Benchmark Research
Previous studies have built benchmark for graph-related tasks (Errica et al., 2019; Hu et al., 2020b). In particular, Errica et al. (2019) is a benchmark including six different GNN models across nine commonly used TUDataset datasets. Open Graph Benchmark (OGB) (Hu et al., 2020b) evaluate differnet graph neural network approaches experiments on graph classification, graph regression, and node classification. Errica et al. (2019) only involve one graph pooling method and Hu et al. (2020b) does not involve any graph pooling methods. In comparison, our method focuses on graph pooling techniques rather than graph neural networks. Moroever, our benchmark explores the robustness of these methods by introducing noise attacks in both graph classification and node classification tasks and investigate their generalizability through out-of-distribution shifts.
Appendix B Details of U-Net based Node Classification
Most existing studies that combine node classification with pooling have utilized the U-Net architecture (Wu et al., 2022a; Song et al., 2024; Zhang et al., 2021). Figure A.1 shows the overview of the U-Net framework. Pooling plays a crucial role in the U-Net structure, as it progressively reduces the graph size in the downsampling path to extract important global features while preserving essential local information. In the upsampling path, pooling facilitates the fusion of features from the downsampling path with those being progressively recovered in the upsampling path through skip connections, thereby assisting the model in accurately performing graph node classification or other tasks. Pooling not only simplifies the graph’s complexity but also provides the model with multi-scale feature representation capabilities.
In the downsampling path, the input feature matrix is first subjected to graph convolution, where the product of the adjacency matrix and the feature matrix, along with the weighted sum of the weight matrix and the bias term, yields the activated feature matrix . Next, a pooling operation is applied, reducing the number of nodes by selecting those with higher scores, thereby transforming the original feature matrix ( and adjacency matrix into a smaller feature matrix and a corresponding adjacency matrix . In the upsampling path, the pooled feature matrix is first upsampled to restore the original number of nodes, generating a new feature matrix . Then, the restored feature matrix is concatenated with the corresponding feature matrix from the downsampling path, forming the merged feature matrix . Finally, the merged feature matrix undergoes another graph convolution, resulting in the output feature matrix .
Appendix C Details of Selected Pooling Methods
C.1 Node Dropping Pooling
C.2 Node Clustering Pooling
Appendix D Detailed Description of Datasets
D.1 Graph Classification
@@
AIDS consists of 2,000 graphs. Each graph corresponds to a molecule, where the nodes represent individual atoms and the edges represent chemical bonds between these atoms. The goal is to predict the inhibitory effect of molecules on HIV based on their structure. (Riesen & Bunke, 2008).
@@
Mutagenicity contains 4,337 molecular graphs. In Mutagenicity, each graph represents a molecule, where the nodes are atoms and the edges denote chemical bonds between the nodes. The classification goal is to predict whether a molecule is mutagenic or not (Debnath et al., 1991).
@@
D.2 Graph Regression
Table A.2 provides an overview of the selected datasets in terms of their tasks, compounds and their features, recommended splits, and metrics. A more detailed description is provided below.
@@ D.3 Node Classification
D.4 Out-of-distribution shifts
Size shifts. For the selected datasets, including NCI1, D&D, NCI109, and IMDB-B, we utilized the data provided by the authors of size-invariant-GNNs (Bevilacqua et al., 2021). In this setup, the graphs with the smallest 50% of nodes are used as the training set, those with the largest 20% of nodes are used as the test set, and the remaining graphs were used as the validation set. The data can be downloaded from https://www.dropbox.com/s/38eg3twe4dd1hbt/data.zip.
@@ D.4 Out-of-distribution shifts
Appendix E Additional Experimental Details
E.1 Graph Classification
@@ E.4 Hyperparameter Tuning
Appendix F Additional Experiments
F.1 Robustness Analysis
@@
F.2 Generalizability Analysis

### v4

@@ 4.3 Robustness Analysis
Table 6 presents the results of the robustness analysis for node-level tasks. From Table 6, we observe the following: Firstly, random attacks on the graph lead to a decrease in performance on node classification tasks, with different types of attacks causing varying degrees of performance degradation. Randomly adding edges has the most negative impact on performance, while randomly deleting edges has the least impact. Secondly, for larger graphs such as Cora, Citeseer, and Pubmed, KMISPool performs the best, whereas for smaller graphs such as Cornell, Texas, and Wisconsin, ASAPool performs better. Thirdly, for graphs with a large diameter, such as PubMed (Diameter=18), SAGPool performs best when edges are deleted or node features are masked. SAGPool leverages graph convolution to compute self-attention scores, enabling node importance to depend on both its intrinsic features and its topological relationships with other nodes. This mechanism allows SAGPool to effectively capture long-range dependencies in large-diameter graphs, as the attention mechanism can focus on important nodes even when long connection paths exist. Appendix C.1 provides results for the robustness analysis of node-level tasks.
As shown in Figure 3, the model’s performance generally declines as the noise intensity increases. It is observed that, at the same level of noise, the impact on accuracy is more pronounced on smaller datasets Cora and CiteSeer, while it is relatively minor on larger dataset Pubmed. Among the three types of noise, although the accuracy of nearly all methods decreases amidst fluctuations, KMISPool and PANPool exhibit the strongest robustness, while COPool performs relatively poorly, despite the fact that most pooling methods show very similar performance.
@@ 4.4 Generalizability Analysis
Table 7 and Table 8 presents the performance of different graph pooling methods under out-of-distribution shifts. For the graph-level datasets D&D and NCI1, we implement two types of distribution shifts. The first type is based on the number of nodes, where the smallest 50% of graphs by node count are used as the training set, and the largest 20% as the test set, with the remainder serving as the validation set (Bevilacqua et al., 2021; Chen et al., 2022). Following the same criteria, the second type of out-of-distribution shifts are generated based on graph density (Chen et al., 2022). For the node-level datasets Cora and Citeseer, the first type of out-of-distribution shift is the top 50% of nodes with the highest degrees as the training set, the bottom 25% with the lowest degrees as the test set, and the remaining nodes as the validation set. The second type is based on closeness centrality (the reciprocal of the sum of the shortest path lengths from a node to all other nodes). We use the 50% of nodes with the lowest closeness as the training set, the 25% with the highest closeness as the test set, and the remaining nodes as the validation set. For further details and more experiments for the generalizability analysis, please refer to the Appendix A.4 and Appendix C.2.
@@ 4.5 Further Analysis
Backbone Analysis. Figure 6 presents the performance of four pooling methods based on GCNConv (Kipf & Welling, 2016), GATConv (Veličković et al., 2017), SAGEConv (Hamilton et al., 2017), and GraphConv (Morris et al., 2019) on four datasets, NCI1, NCI109, PROTEINS, and PROTEINS_FULL. On average, as the backbone models change, most pooling methods exhibit significant performance fluctuations, and no single backbone model consistently maintains a leading position. Except for the PROTEINS_FULL, the performance of GraphConv is relatively better.
@@
5 Related Work
5.1 Graph Classification and Graph Regression
5.2 Node Classification
Node classification aims to classify each node in a graph based on its attributes and relationship with the other nodes (Xiao et al., 2022; Ju et al., 2024; Zhong et al., 2022; Prieto et al., 2023). Node classification has various applications in the real world, including social network analysis (Camacho et al., 2020), knowledge graphs (Ye et al., 2022), bioinformatics (Bhagat et al., 2011) and online commerce services (Yu et al., 2023). Graph neural networks have been widely utilized to solve the problem by learning semantics information across nodes by neighborhood propagation. Since graph pooling would reduce the number of nodes, recent works utilize a U-Net architecture (Gao & Ji, 2019), which involves down-sampling and up-sampling with residue connections. In this work, we systematically evaluate the performance of graph pooling on node classification and observe that graph pooling has limited improvement compared with basic graph neural network architectures.
5.3 Previous Benchmark Research
Previous studies have built benchmarks for graph-related tasks (Errica et al., 2019; Hu et al., 2020b; Tönshoff et al., 2023). In particular, Errica et al. (2019) is a benchmark including six different GNN models across nine commonly used TUDataset datasets. Open Graph Benchmark (OGB) (Hu et al., 2020b) evaluates different graph neural network approaches experiments on graph classification, graph regression, and node classification. Errica et al. (2019) only involve one graph pooling method and Hu et al. (2020b) does not involve any graph pooling methods. In comparison, our method focuses on graph pooling techniques rather than graph neural networks. Moreover, our benchmark explores the robustness of these methods by introducing noise attacks in both graph classification and node classification tasks and investigating their generalizability through out-of-distribution shifts.
6 Summary of Observations and Guidance
Overall, we summarize the main insights basde on our experimental results. Firstly, for graph classification and regression tasks on small-scale graphs with fewer nodes and edges, we recommend using feature extractors that aggregate information from neighboring nodes, such as GCNConv, along with MLPs, to capture both the topological semantics of the graph and the feature semantics of the nodes (e.g., GSAPool and CGIPool). This is because, in small-scale graphs, the topological information of the nodes is relatively easy to obtain. Conversely, for graph classification and regression tasks on large-scale graphs with a substantial number of nodes and edges, we suggest focusing on both local structures and global connectivity patterns (e.g., ParsPool and SEPool), paying attention not only to the topological information of the nodes but also to the link structures of the edges.
Secondly, for node classification tasks on graphs with strong connectivity, we recommend designing pooling structures that aggregate feature information from nodes and their neighbors during the pooling process (e.g., GSAPool and SAGPool). The reason is that in such graphs, nodes are closely connected, and the information between neighboring nodes is richer and easier to extract. On the other hand, for node classification tasks on graphs with weak connectivity, we suggest designing adaptive pooling algorithms that focus more on global information and select nodes more uniformly across the graph (e.g., KMISPool). This is because weaker connectivity implies a more dispersed graph topology, and pooling methods that focus solely on local information may lose representativeness in node selection.
Thirdly, for node classification tasks on graphs with a large diameter, we recommend combining attention mechanisms with algorithms that uniformly select nodes (e.g., SAGPool and KMISPool). The attention mechanism can capture important nodes in graphs with long-range topological dependencies, while uniform node selection allows for the capture of long-distance dependencies and enhances the representativeness of the selected nodes.
Fourthly, when facing graph structural noise attacks, for node classification tasks, we recommend designing pooling algorithms that uniformly sample nodes from the graph. This approach is less sensitive to graph edges and node features compared to algorithms that aggregate local information (e.g., KMISPool). For graph classification tasks, we suggest aggregating neighbor features when designing node dropping pooling methods (e.g., DiffPool, MincutPool, and JustBalancePool). This dilutes the outliers of individual nodes, making the representations of higher-level supernodes more dependent on the overall neighborhood information. When designing node clustering pooling, multiple rounds of neighbor message passing should be performed, as noise features are averaged or weighted multiple times during propagation, enhancing noise suppression. Additionally, when designing the loss function, attention should be paid to balancing the size of each cluster to prevent noise nodes or edges from causing the model to collapse into extreme allocations.
Fifthly, on large-scale graphs, graph pooling can improve computational efficiency and allow graph neural networks to capture more important nodes. Although node clustering pooling shows strong performance (e.g., AsymCheegerCutPool, DiffPool, MincutPool, ParsPool), its computational and spatial costs are higher than those of node dropping pooling (e.g., TopKPool, SAGPool). The reasoning is that the pooling structure requires a soft assignment matrix to determine the clustering relationships of all nodes. Future research could focus on how to efficiently compute this soft assignment matrix.
@@ 7 Conclusion
Appendix A Detailed Description of Datasets
A.1 Graph Classification
@@
AIDS consists of 2,000 graphs. Each graph corresponds to a molecule, where the nodes represent individual atoms and the edges represent chemical bonds between these atoms. Here, we want to predict the inhibitory effect of molecules on HIV based on their structure. (Riesen & Bunke, 2008).
@@
Mutagenicity contains 4,337 molecular graphs. In Mutagenicity, each graph represents a molecule, where the nodes are atoms and the edges denote chemical bonds between the nodes. The classification goal is to predict whether a molecule graph is mutagenic or not (Debnath et al., 1991).
@@
A.2 Graph Regression
Table A.2 provides an overview of the selected datasets. A more detailed description is provided below.
@@ A.3 Node Classification
A.4 Out-of-distribution shifts
Size shifts. For the selected datasets, including NCI1, D&D, NCI109, and IMDB-B, we utilized the data provided by the authors of size-invariant-GNNs (Bevilacqua et al., 2021). In this setup, the graphs with the smallest 50% of nodes are used as the training set, those with the largest 20% of nodes are used as the test set, and the remaining graphs were used as the validation set.
@@ A.4 Out-of-distribution shifts
Appendix B Additional Experimental Details
B.1 Graph Classification
@@ B.4 Hyperparameter Tuning
Appendix C Additional Experiments
C.1 Robustness Analysis
@@
To generate real-world noise due to spurious or missing links, we follow current works (Xia et al., 2023; Zhao et al., 2024) to build a kNN graph instead of randomly adding noise. The performance can be found as in Table LABEL:noise_node_real. From the results, we can find that the differences among various pooling methods are relatively small, which aligns with the observations from noise attacks involving random addition/removal of edges and random masking of node features. Basically, KMISPool, GSAPool, and HGPSLPool demonstrate the best overall performance. We have also introduced the label noise by randomly modifying the class labels of 30% of the nodes. From the results in Table LABEL:noise_node_label, we can observe a similar observation.
| Dataset | TopKPool | SAGPool | ASAPool | PANPool | COPool | CGIPool | KMISPool | GSAPool | HGPSLPool |
| Cora | 74.160.29 | 74.000.59 | 74.520.08 | 74.090.32 | 73.100.22 | 73.670.05 | 74.590.27 | 74.240.25 | 74.730.23 |
| CiteSeer | 72.310.07 | 72.480.10 | 72.680.19 | 72.670.17 | 72.490.16 | 72.310.21 | 72.740.28 | 72.490.09 | 72.610.21 |
| PubMed | 78.650.11 | 78.570.06 | OOM | 77.710.05 | 78.430.17 | 78.350.04 | 78.840.09 | 78.680.09 | OOM |
| Dataset | TopKPool | SAGPool | ASAPool | PANPool | COPool | CGIPool | KMISPool | GSAPool | HGPSLPool |
| Cora | 61.160.11 | 61.160.23 | 61.600.08 | 61.560.27 | 60.650.14 | 60.860.28 | 61.620.16 | 61.210.15 | 61.640.17 |
| CiteSeer | 58.630.67 | 53.250.30 | 53.540.23 | 53.490.13 | 52.760.18 | 52.600.33 | 53.710.22 | 53.370.06 | 53.720.10 |
| PubMed | 62.420.09 | 62.370.08 | 62.500.11 | 62.220.04 | 62.140.08 | 62.020.16 | 62.280.07 | 62.470.08 | 62.160.07 |
C.2 Generalizability Analysis
@@
C.3 Hyperparameter Sensitivity Analysis
Figure A.1 presents the results of a hyperparameter sensitivity analysis conducted on several pooling methods, evaluated on MUTAG and IMDB-MULTI. In particular, we conduct experiments on MUTAG and IMDB-MULTI by varying the number of hidden channels in and learning rates in for the pooling methods. From the results, we can find that the performance of pooling methods fluctuates with changes in hidden channels and learning rates. For smaller datasets such as MUTAG, hidden channels of 32 and 64 combined with smaller learning rates yielded better performance. For larger datasets such as IMDB-MULTI, hidden channels of 256 and slightly larger learning rates demonstrated superior performance.
C.4 Visualization
Figure A.2 visualizes the first and second pooling steps of GSAPool, TopKPool, and SAGPool, as well as the original unpooled graph, on the third graph of the MUTAG dataset. For small-scale graphs such as MUTAG (Avg. nodes=17.9, Avg. edges=39.6), GSAPool partitions the graph into distinct "communities" during pooling, while TopKPool and SAGPool maintain graph connectivity. This demonstrates the superior performance of GSAPool, as it retains nodes from different semantic units (e.g., functional groups), whereas TopKPool and SAGPool tend to select nodes from a single dense region (e.g., the molecular backbone), potentially losing critical functional group information. This further validates the advantage of GSAPool, as its dual-modal filtering of structural information (SBTL) and node feature information (FBTL) accurately identifies community cores, enabling the classifier to directly capture chemical patterns. Figure A.3 and Figure A.4 illustrate the first and second pooling steps of KMISPool, TopKPool, and SAGPool, along with the original unpooled graph, on the Texas and Cora datasets. For datasets with weaker connectivity, such as Texas (Avg. degree=3.2), and those with higher clustering coefficients, such as Cora (Avg. CC=0.24), KMISPool retains more central nodes and preserves connectivity more effectively. In contrast, SAGPool and TopKPool retain fewer nodes and often fail to preserve high-degree nodes. SAGPool and TopKPool rely heavily on local features, such as node degrees or attention scores, for node selection. In weakly connected graphs, this local strategy may lead to the omission of high-degree nodes, as they may be concentrated in specific regions, while other regions are oversampled. In comparison, KMISPool achieves more uniform sampling by leveraging its global node selection strategy.